5. A spool of Cu wire has a starting diam. Of 2.5 mm. It is drawn through a die with an opening of 2.1 mm. Coefficient of friction at the work-die interface is 0.08. The pure Cu has a strength coefficient of 300 MPa & a strain hardening exponent of 0.5. The operation is done at room temperature. Determine: (A) area reduction, (B) draw stress & C) draw force required for the operation. Assume die angle to be 100.
Answer #1

(a) Area Reduction() γ= 0.244 (b Draw Stress ( Ae 3.46 0-300(e) 300(.0 .3(a) -=177, 22 MPa . ivem strain hardeniną expotS D 2
Answer #2

·             A) r=()=()=0,293

B) =ln= =0,348
Yf=
==271Mpa
D=
==2,3
Lc===0,647mm
=0,88+0,12()=1,3
d
=Yf(1+()ln = d=271(1+()1,3*ln=153 MPa
C) f=Af* d
=3,46*153=528,64N


source: YONAS
answered by: YONAS
Answer #3

·             A) r=()=()=0,293

B) =ln= =0,348
Yf=
==271Mpa
D=
==2,3
Lc===0,647mm
=0,88+0,12()=1,3
d
=Yf(1+()ln = d=271(1+()1,3*ln=153 MPa
C) f=Af* d
=3,46*153=528,64N


source: YONAS
answered by: YONAS